OpenStudy (anonymous):

Use a both elimination and substitution to solve this system of linear equations. Type your answer as an ordered triplet, like this: (-2, 4, 13)

OpenStudy (anonymous):

\[2x - 10y + 1 = z \]

OpenStudy (anonymous):

\[4x + y = z + 18\]

OpenStudy (anonymous):

\[-2x - 8z = y - 1 \]